OpenStudy (anonymous):

Solve for y using separation of variables dy/dx= (1+y^2)e^x

OpenStudy (anonymous):

Get all the terms involving x on one side, and all the terms involving y on the other: (1/(1+ y^2)) dy = e^x dx Integrate both sides: arctan(y) = e^x + C take tangent of both sides: y = tan(e^x + C)

OpenStudy (anonymous):

yeppers

OpenStudy (anonymous):

Thanks!